40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 11373     Entry time: Wed Jun 24 08:05:43 2015
Author: Steve 
Type: Update 
Category: PSL 
Subject: laser turned on 

The laser went off around 11am yesterday. It was turned on

Attachment 1: laserWasOff.png  50 kB  Uploaded Wed Jun 24 11:37:18 2015  | Hide | Hide all
laserWasOff.png
ELOG V3.1.3-